Paweł Dybiec

pdybiec@stud.cs.uni.wroc.pl Computer Science and Mathematics Student

HR Department Corporation 123 Pleasant Lane 12345 City, State February 14, 2017

Dear Sir or Madam,

Lore ipum dolor sit amet, consectetuer adiscipliging elit. Up purus ... Sincerely yours,

Paweł Dybiec

Attached: curriculum vitæ

Paweł Dybiec

pdybiec@stud.cs.uni.wroc.pl Computer Science and Mathematics Student

Education

University of Wrocław  Wrocław
BSc  2014-now

Individual Studies in Computer Science and Mathematics (ISIM)

Implementation of both programs (CS and Math) on higher tier courses and groups dedicated for ISIM students

14th Secondary School  Wrocław

2011-2014

Courses of Mathematics, Computer Science at the extended level

Experience

14th Secondary School  Wrocław
teacher for middle school students  2015,2016

Conducting classes for students preparing for Polish Olympiad in Informatics for Middle School Students

14th Secondary School  Wrocław
teacher at the Educational Camp for secondary school students2014,2015,2016

conducting classes for students preparing for Polish Olympiad in Informatics

Plus GPS Q4B  Wrocław
C++ Summer trainee  2015

Skills

Programming languages: C++, Python, OCaml Tools: Linux, git, bash Others: Algorithms, Operating systems, Functional Programming,Statistics, Monte Carlo methods Languages: Fluent in English language

Comptetive Programming

AMPPZ: Polish Collegiate Programming Contest OI: Polish Olympiad in Informatics MWPZ: Local, open ICPC-like contest

Other

https://github.com/dyniec/:
http://codeforces.com/profile/dyniec:

Hobbies: tabletop games, books, comptetive programming